OpenStudy (anonymous):

Find the sum of all values of k so that the graph of y = x^2 - 2x + 7 and the graph of y = kx + 5 are tangent.

OpenStudy (asnaseer):

do you know how to find the slope of the tangent line at any point for the curve: \(y=x^2-2x+7\)?

OpenStudy (anonymous):

no

OpenStudy (asnaseer):

have you studied calculus (specifically differentiation)?

OpenStudy (anonymous):

I haven't

OpenStudy (anonymous):

I was taught basic derivatives though

OpenStudy (asnaseer):

so can you find the derivative of the curve: \(y=x^2-2x+7\)?

OpenStudy (anonymous):

I don't remember much though, but is it 2x - 2

OpenStudy (asnaseer):

that is correct

OpenStudy (asnaseer):

now the derivative represents the slope of the tangent line to the curve \(y=x^2-2x+7\) at any point.

OpenStudy (asnaseer):

so what you need to do is to imagine some point - lets call it \(x=x_1\) at which the slope of the tangent lie to this curve is equal to k.

OpenStudy (asnaseer):

also at that point, both the line \(y=kx+5\) and the curve \(y=x^2-2x+7\) have the same x and y values

OpenStudy (asnaseer):

understand so far?

OpenStudy (anonymous):

can u start from after the derivative part again

OpenStudy (anonymous):

so the derivative of the curve is 2x - 2 then what do u do?

OpenStudy (asnaseer):

let me sketch a graph that will help explain this

OpenStudy (asnaseer):